# Examining the Expanding Role of Synthetic Data Throughout the AI Development Pipeline

### v1

@@ Abstract.
copyright: none
@@ 1. Introduction
Data is a critical building block of AI systems (Halevy et al., 2009). As the tech industry shifts attention and resources towards building large-scale, data-hungry generative models, traditional data sources are being exhausted, giving rise to an intense search for more data (Roose, 2024; Metz et al., 2024). However, developing high-quality datasets remains complex and resource intensive, with substantial costs, time, and expertise required for data collection, annotation, and curation (Sambasivan et al., 2021). Increasingly, synthetic data—artificial data generated using models or algorithmic techniques ((2024), PDPC)—is seen as an appealing alternative (Nikolenko, 2021). OpenAI (Achiam et al., 2023), Apple (Inc., 2024; Gunter et al., 2024), Microsoft (Abdin et al., 2024), Google (Team et al., 2023), Meta (Grattafiori et al., 2024), and IBM (IBM, 2023) have all reported using synthetic data in their AI development pipelines or advertised the ability of their models to create synthetic data.
@@
Now, the widespread availability of large, off-the-shelf generative models has given practitioners the ability to automate the creation of synthetic data without requiring extensive domain-specific expertise or custom-built tools. This has made synthetic data more accessible and rapidly expanded the scope of its use in AI development (Jordon et al., 2022; Liu et al., 2024b). It is increasingly common for researchers and practitioners to use one large generative model (which we refer to as an auxiliary model throughout this paper) to generate synthetic data that is used to train or evaluate another (the primary model). At the same time, the nature of data is evolving. Auxiliary models are used not only to expand training datasets (Luo et al., 2023), create benchmarks (Hartvigsen et al., 2022), and develop test cases (Feng et al., 2023; Perez et al., 2022), but to generate scores or labels for model outputs (Zheng et al., 2023; Dubois et al., 2024), a task previously performed exclusively by human annotators. And since auxiliary models can dynamically generate responses to a primary model’s outputs in real-time, synthetic data no longer needs to be static but can be generated interactively as needed.
@@
One challenge in creating comprehensive guidance on synthetic data is a lack of understanding of practitioners’ current practices. With the landscape rapidly evolving and the use cases so varied, we do not have a clear picture of the full spectrum of practitioners’ motivations for using synthetic data, the properties they hope it will satisfy, or how (or even if) they validate the data they create. To address this gap, we explore the following research questions:
@@
Our interviews reveal how auxiliary models are now embedded in nearly all stages of the AI development pipeline, with synthetic data increasingly substituted for traditional human-generated data and labels. [^1] While we use terms like ’real’ or ’human-generated’ data for distinction throughout this paper, we acknowledge that synthetic data is designed through human intervention, too. Ontological questions about what qualifies as synthetic data are important but they are beyond the scope of this paper. Practitioners perceive synthetic data as a promising and indispensible tool to address data scarcity challenges, resource constraints, and high data collection costs, across a range of critical tasks, including evaluating harms, diversifying datasets to reduce bias, preserving privacy when working with sensitive data, and scoring model outputs. However, practitioners faced challenges in understanding and controlling the outputs of auxiliary models and generating data that accurately depicts members of underrepresented groups. And although practitioners acknowledged the importance of data validation, they struggled to define what makes synthetic data ‘good’ and reported that most validation currently takes the form of ‘spot-checking’ or ‘eyeballing.’ Responsible AI experts noted additional limitations, including risks of stereotyping and the removal of avenues for data subjects to exercise agency over their data.
We reflect on these findings and what they mean for the evolution of data practices in section 5 and conclude by highlighting opportunities for further FAccT research to support the responsible use of synthetic data.
@@ 2. Related Work
In this section, we overview uses of synthetic data in modern AI development pipelines, draw attention to critiques of synthetic data, and engage with scholarship examining industry practices around AI development and data production.
@@ 2.1. Synthetic Data Usage in Modern AI Development
Within the training stage of the AI development pipeline, synthetic data is used to improve generative model capabilities. For instance, researchers have designed synthetic datasets geared to teaching models mathematical reasoning (Luo et al., 2023; Yu et al., 2023) and computer programming (Hu et al., 2023). Synthetic language-image datasets have been designed to improve the performance of multi-modal applications (Liu et al., 2024a), while datasets of synthetic multilingual question-answer pairs have been used to improve cross-lingual performance (Riabi et al., 2020; Shakeri et al., 2020). Synthetic data is also commonly used for knowledge distillation, which aims to transfer knowledge from a large model to a smaller one (Hinton et al., 2015; Gou et al., 2021). Across these varied use cases, synthetic data is positioned as a way to augment real-world datasets and introduce controlled variability (Liu et al., 2024b).
Within the evaluation stage, researchers have generated synthetic test cases to evaluate models for characteristics like factual consistency (Feng et al., 2023) and harmful behavior (Perez et al., 2022), and have published synthetic benchmark datasets, for instance, for detecting harmful output (Hartvigsen et al., 2022). Synthetic test cases can take the form of static model prompts or, since auxiliary models can generate responses on the fly, interactive simulations of a user’s interactions with a model or system. Beyond creating test cases, auxiliary models are increasingly used to evaluate or score outputs from a primary model (Gilardi et al., 2023; Dubois et al., 2024), an idea colloquially known as ‘LLM-as-a-judge’ (Chiang and Lee, 2023). We view this as a form of synthetic data as the scores produced by the auxiliary model take the place of labels that previously would have been collected from human evaluators.
@@ 2.2. Critical Studies of Synthetic Data
Scholars have challenged the assumption that data offers an unmediated reflection of reality, revealing instead its deeply constructed nature. As Gitelman (Gitelman, 2013) argued, data is never ‘raw.’ The abstraction introduced by synthetic data creates further distance from the material realities seemingly represented in datasets, raising the question of whether and to what extent synthetic data is a reflection of reality.
@@ 3. Methods
To further explore the limitations and ethical considerations of the practices identified as well as paths forward, the second phase of interviews focused on understanding responsible AI (RAI) experts’ perspectives on the sociotechnical dimensions of synthetic data use. After asking about participants’ areas of focus within RAI, we asked them to reflect on the ways synthetic data intersects with privacy, fairness, consent, and other RAI concerns, discussing whether synthetic data could help mitigate existing challenges and whether it introduces new risks and trade-offs. To scaffold our discussions, we distilled the use cases of synthetic data mentioned in phase 1 interviews into three hypothetical scenarios. For each scenario, participants reflected on the practical and ethical implications of the practice, including any concerns and trade-offs involved in adopting synthetic data over traditional methods. Refer to Appendix A for study materials.
All interviews were conducted over video conferencing software in English between May and August 2024. Participation was voluntary and informed consent was obtained electronically. Participants were informed of the purpose of the study and were told that they could refuse to answer any questions and ask for the recording to be paused at any time. Each session lasted 40–60 minutes, and participants were compensated with a gift card of \$75 USD. We recorded field notes and video recordings, which were transcribed for analysis. All participant data, including interview transcripts and recordings, was stored in a secure location and was not accessible by anyone outside the research team. The research team checked all included quotes to ensure that no identifying information about participants or their organizations was revealed. Our study design was reviewed and approved by our organization’s Institutional Review Board.
@@
Practitioners RAI Experts
Participant ID contains “P”
(n = 19)
@@
Table 1. Summary of participants’ roles. Participants spanned a total of 14 organizations across the United States.
@@
Limitations. Because of the sensitive nature of our study and the current competitive environment in AI, recruiting practitioners to talk about their data practices was challenging. We were limited in the number of participants we could recruit, the breadth of organizations they represented, and the global diversity of those participants. Most AI practitioners in our study focused on language-based technologies (with three exceptions) and were based in the United States. This may lead to biases or gaps in our findings. Additionally, while part of our original motivation for recruiting RAI experts was to aggregate guidance for synthetic data use, we found that concrete best practices have yet to emerge. We therefore pivoted this discussion to steps the community can take to develop best practices in light of our findings.
@@ 4.1. The Expanding Role of Synthetic Data
Use cases across the AI pipeline. Practitioners described a breadth of use cases for synthetic data, closely aligned with those proposed in the literature (Section 2.1). Within the training stage of the pipeline, practitioners frequently used auxiliary models to produce large datasets or augment existing datasets for pre-training new, primary generative models, particularly in domains where data scarcity or sensitivity limited access to training datasets. Participants explained how synthetic data enabled them to create specialized corpora, such as synthetic legal documents or medical records, which would otherwise remain inaccessible due to privacy and/or compliance requirements. Synthetic data was also used to fine-tune existing primary models for specialized tasks. P15, for example, generated instruction-following datasets to align model outputs with safety requirements, while P12 used a larger auxiliary model as an ‘oracle’ to perform knowledge distillation  (cf. Gou et al., 2021) to train a smaller, more efficient model.
@@
Beyond generating inputs for evaluation, auxiliary models were also used to evaluate the quality of the primary model’s outputs by producing ratings or scores (similar to LLM-as-a-judge  (Zheng et al., 2023)).
In some cases, the scoring prompts and evaluation rubrics themselves were developed using an auxiliary model. Initially, many participants did not consider their use of auxiliary models for scoring model outputs as a form of synthetic data (or ‘data’ at all), but on further reflection, they came to identify these labels as ‘synthetic evaluation data.’
Promises of synthetic data. In describing motivations, practitioners emphasized that auxiliary models enabled them to overcome data and resource constraints and thus significantly accelerate their development and evaluation processes.
@@
The use of human annotators, meanwhile, was limited by the growing needs across teams, especially for evaluation. P5 and P11 noted that tight evaluation timelines and organizational demands required generating thousands of prompts and testing models within just a few weeks, with dozens of teams in P5’s organization conducting weekly model evaluations. They noted that scoring large volumes of model outputs weekly was infeasible with human reviewers.
@@ 4.2. Generating Synthetic Data
Desiderata for synthetic data. The two desiderata most commonly identified by practitioners were diversity and resemblance to ‘natural’ or ‘real’ data. Diversity was often described in terms of capturing a range of scenarios, input types, or demographic characteristics, which participants considered important to enable primary models to generalize effectively across varied real-world contexts. Natural-looking data was described as closely resembling human-generated content, but without replicating the training data of the auxiliary model. For instance, in the context of generating data for automated red-teaming, P3 explained that the average user would not make an explicitly harmful query like “how to make a bomb.” Instead, more frequent “naturally occurring queries” might trigger more subtle forms of harm relating to, for instance, interpersonal bias in the workplace.
Approaches to data generation. Practitioners chose different approaches to synthetic data generation, with distinct trade-offs between control and flexibility. Some preferred a constrained approach that limited the scope of generation to maintain more oversight over outputs. Most commonly, this involved having the auxiliary model generate variations (e.g., paraphrases) of manually curated seed examples.
P13, for example, used the approach of generating synonyms for a list of successful red-teaming prompts, arguing that this structured approach supported the rigor required for controlled experimentation by allowing them to monitor how language shifts impacted the primary model’s responses. Another common constrained approach was to design structured templates that dictate the format the auxiliary model’s output should take, for instance, steering the model output toward specific topics or demographic groups.
Other practitioners were willing to sacrifice some control for the flexibility afforded by more open-ended approaches. One such approach was to generate synthetic data by having the auxiliary model take on a simulated persona reflecting specific demographics or scenarios. Practitioners reasoned that this approach allowed for exploration across a broader set of possible outputs, offering more diversity and resemblance to natural data. While there was a risk of introducing lower-quality irrelevant data, practitioners adopting such open-ended approaches used the strategy of iterating over the prompt for the auxiliary model if the generated data did not match their expectations.
@@
Choosing an auxiliary model. Participants described an ad-hoc approach to selecting which auxiliary model to use for data generation, with choices driven by immediate needs and availability rather than systematic criteria. Many relied on models readily accessible within their organizations or those considered “state-of-the-art.” P9 explained “we’ve always operated on the idea that the newest model is going to be the best and the most helpful for our customers.” P4 described choosing “the most powerful LLM that [their team] feel is economically reasonable to prompt.” Teams often matched models to task complexity, as P18 noted, using GPT-3.5 for “straightforward tasks, such as cleaning up typos or translating text,” while reserving GPT-4 quotas for more complex activities, such as generating math problems.
@@
Typical large, off-the-shelf models undergo alignment processes intended to ensure helpful and safe responses (cf. Ouyang et al., 2022). By design, this would limit the model’s capacity to generate responses containing stereotypes or potentially offensive content. P5 recalled trying to use an ‘aligned’ auxiliary model for harms testing. When the primary model output “I’m sorry, I cannot generate stereotypes,” the auxiliary model would affirm with responses like “Great, you should not,” derailing the evaluation. For this reason, when working on safety-related tasks, such as fine-tuning a primary model to generate safe responses or evaluating for harms, participants often used what they believed were ‘unaligned’ auxiliary models.
@@
Challenges with synthetic data generation. Participants identified several challenges in generating high-quality synthetic data. Most could be traced to the limitations of the auxiliary models themselves since, as P16 emphasized, synthetic data inevitably reflects the capabilities and limitations of the auxiliary model. Participants described auxiliary models as brittle and highly sensitive to minor changes in prompts. P6 referred to prompt engineering as ”more of an art than a science,” noting that small adjustments (e.g., reordering words, changing punctuation, or rephrasing sentences) would significantly impact quality. Practitioners found themselves engaged in constant experimentation, attempting to (prompt) “engineer away” odd behaviors or inconsistencies.
@@
While natural-looking data was desirable, some practitioners questioned whether synthetic data distributions could closely resemble real data. Many were skeptical,
and P7 cautioned that “one would hope that the distributions are quite similar, but they’re not going to be exactly the same on the real and synthetic data.
@@
While controllability was viewed as a major benefit of synthetic data, practitioners often struggled to understand or control the outputs of the auxiliary models. P12, who used an auxiliary model for model distillation, acknowledged how its opacity left their team uncertain about whether the primary model was learning the correct function, noting that “it is hard to say exactly what’s going on here.” P5 encountered a similar issue when using an auxiliary model to create socio-cultural identity prompts to seed safety testing for their primary model. They intended the simulator to generate responses reflecting specific stereotypes for testing purposes but discovered that “all the responses were about tacos.” Upon manual inspection, their debugging revealed how “it was an issue in the pipeline where a single few-shot example that happened to mention Mexican food and tacos was being injected and overriding some other stuff.” And P13, the practitioner who had generated synonyms for a list of successful red-teaming prompts, acknowledged that even with this targeted method, they couldn’t fully control the model’s semantic interpretations: “If you just think of the word ‘red,’ in addition to giving us colors, [the auxiliary model] gave us this more conceptual kind of bloody. It could’ve gone in a political direction with that. It could’ve gone in a racist direction with that… That might be in the [data], and we currently don’t know.

### v2

@@ Abstract.
journalyear: 2025conference: ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732005isbn: 979-8-4007-1482-5/25/06journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732005isbn: 979-8-4007-1482-5/2025/06
@@ 1. Introduction
Data is a critical building block of AI systems (Halevy et al., 2009). As the tech industry shifts attention and resources towards building large-scale, data-hungry generative models, traditional data sources are being exhausted, giving rise to an intense search for more data (Roose, 2024; Metz et al., 2024). However, developing high-quality datasets remains complex and resource-intensive, with substantial costs, time, and expertise required for data collection, annotation, and curation (Sambasivan et al., 2021). Increasingly, synthetic data—artificial data generated using models or algorithmic techniques ((2024), PDPC)—is seen as an appealing alternative (Nikolenko, 2021). OpenAI (Achiam et al., 2023), Apple (Inc., 2024; Gunter et al., 2024), Microsoft (Abdin et al., 2024), Google (Team et al., 2023), Meta (Grattafiori et al., 2024), and IBM (IBM, 2023) have all reported using synthetic data in their AI development pipelines or advertised the ability of their models to create synthetic data.
@@
Now, the widespread availability of large, off-the-shelf generative models has given practitioners the ability to automate the creation of synthetic data without requiring extensive domain-specific expertise or custom-built tools. This has made synthetic data more accessible and rapidly expanded the scope of its use in AI development (Jordon et al., 2022; Liu et al., 2024b). It is increasingly common for researchers and practitioners to use one large generative model (which we refer to as an auxiliary model throughout this paper) to generate synthetic data that is used to train or evaluate another (the primary model). At the same time, the nature of data is evolving. Auxiliary models are used not only to expand training datasets (Luo et al., 2023), create benchmarks (Hartvigsen et al., 2022), and develop test cases (Feng et al., 2023; Perez et al., 2022), but also to generate scores or labels for model outputs (Zheng et al., 2023; Dubois et al., 2024), a task previously performed by human annotators or with simple automated approaches. And since auxiliary models can dynamically generate responses to a primary model’s outputs in real-time, synthetic data no longer needs to be static but can be generated interactively as needed.
@@
One challenge in creating comprehensive guidance on synthetic data is a lack of understanding of practitioners’ current workflows. With the landscape rapidly evolving and the use cases so varied, we do not have a clear picture of the full spectrum of practitioners’ motivations for using synthetic data, the properties they hope it will satisfy, or how (or even if) they validate the data they create. To address this gap, we explore the following research questions:
@@
While synthetic data is a broad concept, in order to focus on new and evolving practices, we scope our study to the use of data (including labels and scores of outputs) produced by auxiliary generative models for use in the AI development pipeline. While we use terms like ‘real’ or ‘human-generated’ data for distinction throughout this paper, we acknowledge that synthetic data is also shaped by human design and intervention (Lee et al., 2025; Jacobsen, 2023) and that, in practice, there is a spectrum between ‘real’ and ‘synthetic’ data.
Our interviews reveal how auxiliary models are now embedded in nearly all stages of the AI development pipeline, with synthetic data increasingly substituted for traditional human-generated data and labels. Practitioners perceive synthetic data as a promising and indispensable tool to address data scarcity challenges, resource constraints, and high data collection costs across a range of critical tasks, including evaluating harms, diversifying datasets to reduce bias, preserving privacy when working with sensitive data, and scoring model outputs. However, practitioners faced challenges in understanding and controlling the outputs of auxiliary models and generating data that accurately depicts members of underrepresented groups. Although practitioners acknowledged the importance of data validation, they struggled to define what makes synthetic data ‘good’ and reported that most validation currently takes the form of ‘spot-checking’ or ‘eyeballing.’ Responsible AI experts noted additional limitations, including risks of stereotyping and the removal of avenues for data subjects to exercise agency over their data. We reflect on these findings and their implications for the evolution of data practices in section 5 and conclude by highlighting opportunities for future FAccT research to support the responsible use of synthetic data.
Our paper makes three main contributions:
  • We provide an empirical account of the ubiquitous and inconsistent integration of synthetic data into modern AI development pipelines.
  • We highlight limitations and ethical considerations of synthetic data use and propose avenues to study its impact.
  • We provide considerations for the responsible use of synthetic data aimed at both practitioners and policymakers.
@@ 2. Related Work
In this section, we provide an overview of the uses of synthetic data in modern AI development pipelines, draw attention to critiques of synthetic data, and engage with scholarship examining industry practices around AI development and data production.
@@ 2.1. Synthetic Data Usage in Modern AI Development
Within the training stage of the AI pipeline, synthetic data is used to improve generative model capabilities. For instance, researchers have designed synthetic datasets geared to teaching models mathematical reasoning (Luo et al., 2023; Yu et al., 2023) and computer programming (Hu et al., 2023). Synthetic language-image datasets have been designed to improve the performance of multi-modal applications (Liu et al., 2024a), while datasets of synthetic multilingual question-answer pairs have been used to improve cross-lingual performance (Riabi et al., 2020; Shakeri et al., 2020). Synthetic data is also commonly used for knowledge distillation, which aims to transfer knowledge from a large model to a smaller one (Hinton et al., 2015; Gou et al., 2021). Across these varied use cases, synthetic data is positioned as a way to augment real-world datasets and introduce controlled variability (Liu et al., 2024b).
Within the evaluation stage, researchers have generated synthetic test cases to evaluate models for characteristics like factual consistency (Feng et al., 2023) and harmful behavior (Perez et al., 2022), and have published synthetic benchmark datasets, for instance, for detecting harmful output (Hartvigsen et al., 2022). Synthetic test cases can take the form of static model prompts or, since auxiliary models can generate responses on the fly, interactive simulations of a user’s interactions with a model or system. Beyond creating test cases, auxiliary models are increasingly used to evaluate or score outputs from a primary model (Gilardi et al., 2023; Dubois et al., 2024), an idea colloquially known as ‘LLM-as-a-judge’ (Chiang and Lee, 2023). We view this as a form of synthetic data as the scores produced by the auxiliary model take the place of labels that often would have been collected from human evaluators.
@@ 2.2. Critical Studies of Synthetic Data
Scholars have challenged the assumption that data offers an unmediated reflection of reality, revealing instead its deeply constructed nature. As Gitelman (Gitelman, 2013) argued, data is never ‘raw.’
The abstraction introduced by synthetic data creates further distance from the material realities seemingly represented in datasets, raising the question of whether and to what extent synthetic data is a reflection of reality.
@@ 3. Methods
To further explore the limitations and ethical considerations of the practices identified as well as paths forward, the second phase of interviews focused on understanding responsible AI (RAI) experts’ perspectives on the sociotechnical dimensions of synthetic data use. After asking about participants’ areas of focus within RAI, we asked them to reflect on the ways synthetic data intersects with privacy, fairness, consent, and other RAI concerns, discussing whether synthetic data could help mitigate existing challenges and whether it introduces new risks and trade-offs. To scaffold our discussions, we distilled the use cases of synthetic data mentioned in phase 1 interviews into three hypothetical scenarios: using synthetic data to fine-tune models, create evaluation datasets, or simulate user interactions. For each scenario, participants reflected on the practical and ethical implications of the practice, including any concerns and trade-offs involved in adopting synthetic data over traditional methods. Refer to Appendix A for study materials, including the vignettes.
All interviews were conducted over video conferencing software in English between May and August 2024. Each session lasted 40–60 minutes, and participants were compensated with a gift card of \$75 USD. We recorded field notes and video recordings, which were transcribed for analysis. See Section 7 for a discussion of research ethics considerations.
@@
Practitioners (n = 19) RAI Experts (n = 10)
@@
Table 1. Summary of participants’ roles. Participants spanned a total of 14 organizations across the United States. In reporting our findings, we use identifiers beginning with ‘P’ for practitioners and ‘E’ for responsible AI experts.
@@
Limitations. Because of the sensitive nature of our study and the current competitive environment in AI, recruiting practitioners to discuss their data practices was challenging. We were limited in the number of participants we could recruit, the breadth of organizations they represented, and the global diversity of those participants. In particular, all practitioners were based in the United States. Additionally, most practitioners in our study focused on language-based technologies and, therefore, text data. The three exceptions to this focused on multimodal data (text and images), image data, and biological sequence data, respectively. These factors potentially skewed the data practices we observed and challenges highlighted in our findings. Additionally, while part of our original motivation for recruiting RAI experts was to aggregate guidance for synthetic data use, we found that concrete best practices have yet to emerge. We therefore pivoted this discussion to steps the community can take to develop best practices in light of our findings.
@@ 4.1. The Expanding Role of Synthetic Data
4.1.1. Use Cases Across the AI Pipeline.
Practitioners described a breadth of use cases for synthetic data, closely aligned with those proposed in the literature (Section 2.1). Within the training stage of the pipeline, practitioners frequently used auxiliary models to produce large datasets or augment existing datasets for pre-training new primary generative models, particularly in domains where data scarcity or sensitivity limited access to training datasets. Participants explained how synthetic data enabled them to create specialized corpora, such as synthetic legal documents or medical records, which would otherwise remain inaccessible due to privacy and/or compliance requirements. Synthetic data was also used to fine-tune existing primary models for specialized tasks. P15, for example, generated instruction-following datasets to align model outputs with safety requirements, while P12 used a larger auxiliary model as an ‘oracle’ to perform knowledge distillation  (cf. Gou et al., 2021) to train a smaller, more efficient model.
@@
Beyond generating inputs for evaluation, auxiliary models were also increasingly used to evaluate the quality of the primary model’s outputs by producing ratings or scores (as in LLM-as-a-judge  (Zheng et al., 2023)), a task that was previously typically carried out by human annotators or more conventional automated metrics (e.g., (Papineni et al., 2002; Fabbri et al., 2021)). For example, P11 was using an auxiliary model for red-teaming purposes: specifically, to evaluate whether the primary model was producing harmful outputs. They explained how the model “output obviously needs to be categorized, and that is done using [auxiliary] models. So we have kind of like scorers […] across different harm categories that sort the outputs, which I think could count as synthetic data insofar as it’s making a judgment on the output of the [primary model].” In some cases, the scoring prompts and evaluation rubrics themselves were developed using an auxiliary model. Initially, several participants did not consider their use of auxiliary models for scoring model outputs as a form of synthetic data (or ‘data’ at all), but on further reflection, they came to identify these labels as ‘synthetic evaluation data.’
4.1.2. Perceived Promises of Synthetic Data.
Practitioners brought up a variety of reasons why they expected synthetic data to be beneficial, which we describe here. However, as we explore in Sections 4.2.3 and 4.3.2 when reflecting on the challenges of generating and validating synthetic data, these promised benefits were not always realized in practice.
In describing motivations for their use of synthetic data, practitioners emphasized that auxiliary models enabled them to overcome data and resource constraints and thus significantly accelerate their development and evaluation processes.
@@
The use of human annotators, meanwhile, was limited by the growing needs across teams, especially for evaluation. P5 and P11 noted that tight timelines and organizational demands required generating thousands of prompts and testing models within just a few weeks, with dozens of teams in P5’s organization conducting weekly model evaluations. They noted that scoring large volumes of model outputs weekly was infeasible with human reviewers.
@@ 4.2. Generating Synthetic Data
4.2.1. Desiderata for Synthetic Data.
The two desiderata most commonly identified by practitioners were diversity and resemblance to ‘natural’ or ‘real’ data. Diversity was often described in terms of capturing a range of scenarios, input types, or demographic characteristics, which participants considered important to enable primary models to generalize effectively across varied real-world contexts. Natural-looking data was described as closely resembling human-generated content, but without replicating the training data of the auxiliary model. For instance, in the context of generating data for automated red-teaming, P3 explained that the average user would not make an explicitly harmful query like “how to make a bomb.” Instead, more frequent “naturally occurring queries” might trigger more subtle forms of harm relating to, for instance, interpersonal bias in the workplace.
4.2.2. Approaches to Data Generation.
Practitioners chose different approaches to synthetic data generation, with distinct trade-offs between control and flexibility. Some preferred a constrained approach that limited the scope of generation to maintain more oversight over outputs. Most commonly, this involved having the auxiliary model generate variations (e.g., paraphrases) of manually curated seed examples. P13, for example, generated synonyms for a list of successful red-teaming prompts, arguing that this structured approach supported the rigor required for controlled experimentation by allowing them to monitor how language shifts impacted the primary model’s responses. Another common constrained approach was to design structured templates that dictate the format the auxiliary model’s output should take, for instance, steering the model output toward specific topics or demographic groups.
Other practitioners were willing to sacrifice some control for the flexibility afforded by more open-ended approaches. One such approach was to generate synthetic data by having the auxiliary model take on a simulated persona reflecting specific demographics or scenarios. Practitioners reasoned that this approach allowed for exploration across a broader set of possible outputs, offering more diversity and resemblance to natural data. While there was a risk of introducing lower-quality irrelevant data, practitioners adopting open-ended approaches would iteratively rewrite the prompt for the auxiliary model if the generated data did not match their expectations.
@@
Participants described an ad-hoc approach to selecting which auxiliary model to use for data generation, with choices driven by immediate needs and availability rather than systematic criteria. Many relied on models readily accessible within their organizations or those considered “state-of-the-art.” P9 explained “we’ve always operated on the idea that the newest model is going to be the best and the most helpful for our customers.” P4 described choosing “the most powerful LLM that [their team] feels is economically reasonable to prompt.” Teams often matched models to task complexity, as P18 noted, using GPT-3.5 for “straightforward tasks, such as cleaning up typos or translating text,” while reserving GPT-4 quotas for more complex activities, such as generating math problems.
@@
Typical large, off-the-shelf models undergo alignment processes intended to ensure helpful and safe responses (cf. Ouyang et al., 2022). By design, this would limit the model’s capacity to generate responses containing stereotypes or potentially offensive content. P5 recalled trying to use an ‘aligned’ auxiliary model for safety testing. When the primary model outputted “I’m sorry, I cannot generate stereotypes,” the auxiliary model would affirm with responses like “Great, you should not,” derailing the evaluation. For this reason, when working on safety-related tasks, such as fine-tuning a primary model to generate safe responses or evaluating for harms, participants often used what they believed were ‘unaligned’ auxiliary models.
4.2.3. Challenges with Synthetic Data Generation.
@@
Participants identified several challenges in generating high-quality synthetic data. Most could be traced to the limitations of the auxiliary models themselves since, as P16 emphasized, synthetic data inevitably reflects the capabilities and limitations of the auxiliary model. Participants described auxiliary models as brittle and highly sensitive to minor changes in prompts. P6 referred to prompt engineering as ”more of an art than a science,” noting that small adjustments (e.g., reordering words, changing punctuation, or rephrasing sentences) would significantly impact quality. Practitioners found themselves engaged in constant experimentation, attempting to (prompt) “engineer away” odd behaviors or inconsistencies.
@@
While natural-looking data was desirable, some practitioners questioned whether synthetic data distributions could closely resemble real data. Many were skeptical, and P7 cautioned that “one would hope that the distributions are quite similar, but they’re not going to be exactly the same on the real and synthetic data.
@@
While controllability was viewed as a major benefit of synthetic data, practitioners often struggled to understand or control the outputs of the auxiliary models. P12, who used an auxiliary model for model distillation, acknowledged how its opacity left their team uncertain about whether the primary model was learning the correct function, noting that “it is hard to say exactly what’s going on here.” P5 encountered a similar issue when using an auxiliary model to create socio-cultural identity prompts to seed safety testing for their primary model. They intended the simulator to generate responses reflecting specific stereotypes for testing purposes but discovered that “all the responses were about tacos.” Upon manual inspection, their debugging revealed how “it was an issue in the pipeline where a single few-shot example that happened to mention Mexican food and tacos [were] being injected and overriding some other stuff.” And P13, the practitioner who had generated synonyms for a list of successful red-teaming prompts, acknowledged that even with this targeted method, they couldn’t fully control the model’s semantic interpretations: “If you just think of the word ‘red,’ in addition to giving us colors, [the auxiliary model] gave us this more conceptual kind of bloody. It could’ve gone in a political direction with that. It could’ve gone in a racist direction with that… That might be in the [data], and we currently don’t know.